# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v2

@@ 3.4 Self-Routing RAG: Inference
As shown in Figures˜1 and 2, SR-RAG inference unfolds in a single left-to-right pass through three steps: source selection, knowledge collection, and answer generation.
@@
A common approach to selecting a knowledge source is to compare the likelihood $p_{M}(\text{{\textless s\textgreater}}|q)$ for each $s\in\mathcal{S}$, against a fixed threshold (Asai et al., 2024; Wu et al., 2024a). However, this approach fails to account for shifts in the LLM’s ability after fine-tuning and lacks fine-grained control over the decision boundary. To make source selection more robust, we propose a dynamic nearest neighbor-based source selection mechanism that builds a policy datastore based on rollouts after the fine-tuning. Concretely, the fine-tuned LLM is first evaluated on a set of question-answer pairs [^2] Empirically, we reuse the training set so that no additional data or supervision is required.. For each pair, the probabilities of generating the answer $a$ conditioned on different knowledge sources are compared to decide the preferred source. We then build a policy datastore mapping each query to its preferred source, using the hidden representation at <EOQ> as the key. At inference, we retrieve $k$ nearest neighbors from the policy datastore and use their source labels to form a distribution over the sources $p_{D}(\text{{\textless s\textgreater}}|q)$. Finally, to select the best source $s\in\mathcal{S}$, we apply a threshold on the product:
$p_{M}(\text{{\textless s\textgreater}}|q)\times p_{D}(\text{{\textless s\textgreater}}|q).$ (7)
While tackling the challenges of source selection due to shifts in the LLM’s ability, this approach also exhibits better interpretability. Since the policy datastore consists of explicit source assignments, it can be audited, modified, and expanded by human experts to steer SR-RAG’s retrieval behavior in a fine-grained manner.
Subsequently, the knowledge from the corresponding source is gathered. If the LLM prefers $S_{i}$, we use greedy decoding to directly verbalize a single knowledge context. After SR-RAG fine-tuning, the generated context serves as a compressed yet high-quality articulation of the parametric knowledge, which would otherwise require compute-expensive knowledge verbalization to elicit. If an external source <s> is selected instead, we pause the decoding and retrieve from knowledge source $s$. With the verbalized or retrieved knowledge appended to the context, the LLM proceeds to generate the final response.
@@ 4.1 Implementation Details of SR-RAG
The main experiments are performed on two knowledge sources [^3] We present multi-source extensions of SR-RAG in Appendix C.: the 2018 English Wikipedia (<Wiki>) as the external knowledge source, and the LLM itself (<Self>) as the internal knowledge source. We use the official Wikipedia embeddings released by Karpukhin et al. (2020) and retrieve at a granularity of 100-word chunks. GenRead (Yu et al., 2023) is used to verbalize diverse knowledge contexts. GenRead clusters zero-shot knowledge verbalizations in as in-context demonstrations to verbalize diverse knowledge. We limit the verbalized knowledge chunks to a maximum of 150 tokens. From each knowledge source, we collect $n$ = 5 knowledge contexts.
@@
We fine-tune on a mixture of six short- and long-form knowledge-intensive datasets: Wizard of Wikipedia (Dinan et al., 2019), Natural Questions (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), OpenBookQA (Mihaylov et al., 2018), ARC-Easy (Bhakthavatsalam et al., 2021), and ASQA (Stelmakh et al., 2022). This mixture of 53,042 instances is a subset of the RAG instruction tuning data shown effective in Asai et al. (2024). After running the data construction algorithm, 46.9% of the instances are labeled with <Self> and the rest are labeled with <Wiki> as the preferred knowledge source, averaged across three LLMs. Full training details are provided in Section˜B.1.
@@ 4.2 Evaluation
We evaluate on a diverse set of four knowledge-intensive NLP tasks. PopQA (Mallen et al., 2023) is a free-formed long-tail open-domain question answering dataset. Following Asai et al. (2024), we use the subset of 1,399 questions that aims to test knowledge of long-tail entities. TriviaQA (Joshi et al., 2017) is an established open-domain question answering dataset that features relatively complex and diverse questions. We use the same test split and retrieval setup as in Asai et al. (2024). PubHealth (Zhang et al., 2023) is a fact-checking dataset focusing on checking claims in the public health domain. Finally, ARC Challenge (Bhakthavatsalam et al., 2021) is a multiple-choice question answering dataset featuring grade-school level science questions. Following common practice, we perform lexical post-process of the model’s output and report accuracy for PubHealth and ARC and substring matching for PopQA and TriviaQA.
@@
We compare SR-RAG with baselines that cover various training and inference strategies. (1) First, using the LLM before fine-tuning, we compare with either always retrieving or always verbalizing with GenRead. (2) As illustrated in Figure˜2, our main baseline is the state-of-the-art prior selective retrieval pipeline, combining the advantage of He et al. (2021), Asai et al. (2024), and Wu et al. (2024a). Specifically, the likelihoods of the LLM generating the answer with and without retrieval are used to create the knowledge source selection label. Then, we fine-tune the LLM for knowledge source selection (among $S_{e}$ and $\phi$) and generate the answer with optional retrieval. At inference, we apply a uniform threshold of 0.2 to the likelihood of the retrieval token for selective retrieval. (3) Always retrieving with the fine-tuned SR-RAG LLM.
@@ 5.1 Knowledge Verbalization Alters Knowledge Source Preference
To motivate our approach, we first show that knowledge verbalization substantially impacts on identifying when an LLM requires retrieval. Figure˜3 presents a pilot study on Llama-2-7B-Chat using four datasets from SR-RAG’s training mixture. For each instance, we compute the likelihood of the LLM generating the correct answer with no context (blue), with the most helpful GenRead verbalization ($c_{i+}$, orange), and with the most helpful retrieved passage ($c_{e+}$, red). Notably, GenRead changes the preferred knowledge source in 16% of Natural Questions examples and over 30% in the other datasets. This suggests that prior selective retrieval methods, which omit verbalization, may significantly underestimate the LLM’s capabilities, reinforcing the need to incorporate knowledge verbalization for accurate source preference labeling.
@@
Table 2: Source selection accuracy measured on Llama-2-7B-Chat. SR-RAG achieves the best averaged performance in both evaluation settings.
Figure 4: Accuracy and system latency of SR-RAG fine-tuned Llama-2-7B-Chat with different verbalization frequencies. SR-RAG’s source selection policy (marked with stars) achieves near-optimal accuracy-efficiency trade-off without dataset-specific thresholds.
@@ 5.2 SR-RAG Improves Accuracy with Reduced Retrieval Frequency
Table˜1 shows the end-to-end generation performance on three LLMs, demonstrating the advantage of SR-RAG over always retrieving, selective retrieval, and other baselines. Relative to always retrieving with the pretrained model, SR-RAG improves accuracy by 19.3% (Llama-2-7B-Chat), 13.4% (Phi-3.5-Mini-Instruct), and 5.9% (Qwen2.5-7B-Instruct). Against the strongest selective retrieval baseline, SR-RAG yields 8.5%/2.1%/4.7% higher accuracy. Interestingly, despite saving 8–15% retrievals, the selective retrieval baseline achieves nearly identical accuracy to always retrieving. This suggests that selective RAG alone cannot reliably identify queries that do not require retrieval. In contrast, SR-RAG reduces 20–40% retrievals while maintaining or improving performance thanks to accurate knowledge source selection and high-quality verbalization.
Remarkably, with a uniform inference datastore and threshold, SR-RAG dynamically adapts its retrieval behavior based on dataset characteristics. For a dataset that emphasizes long-tail knowledge like PopQA, SR-RAG tends to retrieve external knowledge most of the time. On the other hand, for PubHealth and ARC where the model’s knowledge may suffice for a number of questions, SR-RAG relies on internal knowledge more confidently, resulting in much better performance compared to always retrieving. Qualitative examples illustrating SR-RAG’s behavior can be found in Figure˜7 and Figure˜8 in the appendix.
@@ 5.3 SR-RAG Makes Accurate Source Selection Decisions
Can SR-RAG accurately choose the correct knowledge source? Table˜2 compares SR-RAG with Self-RAG (Asai et al., 2024) and an SR-RAG variant without kNN-based policy. We evaluate accuracy using two criteria: (1) abstaining from retrieval when it does not harm performance (top) and (2) retrieving only when it provides strictly better context (bottom). While all methods perform well under the first criterion, SR-RAG achieves the highest overall accuracy. Under the second, stricter definition, SR-RAG demonstrates substantially better performance, outperforming Self-RAG by 14.7% in average AUROC. Removing kNN-based policy (w/o. kNN) leads to large drops in both accuracy and AUROC, highlighting the benefit of adapting to fine-tuning-induced shifts in LLM ability.
@@ 5.4 System Efficiency
We assess SR-RAG’s end-to-end latency using Llama-2-7B-Chat under batched inference [^6] Detailed latency modeling formulation in Section B.2. In our setting, latency is linearly correlated with retrieval frequency under small batch sizes.. Figure˜4 shows the trade-off between accuracy and latency under varying retrieval proportions. As expected, latency improves as fewer instances invoke retrieval. Due to dataset-specific difficulty, the optimal retrieval proportion varies per dataset. However, SR-RAG’s learned source selection policy achieves a near-optimal accuracy-efficiency trade-off across all datasets, without any dataset-specific threshold tuning. This confirms that SR-RAG not only improves performance but does so efficiently.
@@ 5.5 Ablation Study
To understand the contribution of each component, we perform ablations on Llama-2-7B-Chat (Table˜3). Removing kNN-based source selection (w/o. kNN) results in lower accuracy and higher retrieval rates, confirming that kNN inference helps the model adapt to ability shifts introduced by fine-tuning. Disabling knowledge verbalization during preference labeling (w/o. kv. label) causes the LLM to over-rely on retrieval and reduces performance, showing that verbalization is crucial for effective labeling. Finally, $\mathcal{L}_{stage2}$ is ablated and the $\mathcal{L}_{stage1}$ loss is kept for stage 2 training. The result shows significantly more retrievals as the model’s knowledge verbalization ability degrades. These findings highlight that all components are vital to SR-RAG’s success. In Appendix˜D, we further analyze the <EOQ> representations, other SR-RAG hyperparameters, as well as qualitative examples.
@@ 6 Conclusion
Appendix A List of Notations
@@
Table 4: A summary of the key symbols and parameters used in the paper.
Appendix B SR-RAG: Further Details
B.1 Training Details
@@
We implement GenRead following the setup in Yu et al. (2023). In the second verbalization round, five clusters with five in-context examples each are used. All datasets except ASQA follow the general prompt shown in Figure˜5. For ASQA, we include an additional instruction to handle ambiguity: “If the question is ambiguous, generate multiple documents for each possibility.
@@
During training, the loss is computed on the knowledge part only if <s> is <Self>. If <s> is <Wiki>, we augment $c_{e+}$ by randomly appending $\max(p-1,0)$ retrieved contexts, where $p$ is sampled from a Poisson distribution with $\lambda=2$. This data augmentation improves the LLM’s robustness to different retrieval strategies and various levels of retrieval quality. For stage 1 training, we use batch size 64, learning rate 1e-5, and fine-tune for 1 epoch. For stage 2 training, we use batch size 64, learning rate 5e-7, $\beta$ = 0.3 for DPO, and train for another epoch. All the experiments are performed on a machine with eight A800 (80GB) GPUs and a machine with eight A6000 GPUs. On eight A800 (80GB) GPUs, the two-staged training takes approximately 10 hours for a 7B-sized model.
B.2 Latency Formulation
@@
  • Verbalization Latency ($T_{v}$): The time for the LLM to verbalize parametric knowledge if itself is selected as the knowledge source.
@@
  • As the nearest neighbor only involves one matrix product for similarity calculation and one top-k operation, $T_{d}$ is generally very small. In fact, we find $T_{d}$ (0.01s) <<<< $T_{g}$ (0.1s) << $T_{v}$ (1s). On the other hand, $T_{rs}$ is the major bottleneck of the pipeline. As a result, in an online setting, the system’s efficiency gain directly converges to the percentage of retrievals it is able to avoid.
Appendix C SR-RAG: Multi-Source Extensions
C.1 Experiment: Routing between General and Specialized Sources
We perform experiments on the PubMed corpus (National Center for Biotechnology Information (), NCBI), which contains professional biomedical literature. Compared to Wikipedia and the LLM itself, PubMed has the unique advantage of featuring long-tail specialized biomedical knowledge. Meanwhile, it falls short at providing knowledge for many general factuality questions, posing a challenge to accurately routing the knowledge source.
Based on this new setup, we formulate SR-RAG with three knowledge source tokens: <Self>, <Wiki>, and <Pubmed>. We leverage the MedRAG toolkit (Xiong et al., 2024) to implement retrieval on PubMed. The training and inference strategies of the two-source SR-RAG recipe naturally extends to this formulation and are followed in our experiments here. For training and evaluation, we additionally use PubMedQA (Jin et al., 2019), a widely used biomedical question answering dataset. PubMedQA features short questions extracted from PubMed abstracts and provides associated ground truth yes/no/maybe answers. We sample 10k artifically generated documents for training (combined with the data mixture used in the main text) and 500 examples from the expert labeled subset for evaluation. The same set of LLMs from three families (Llama, Qwen, Phi) are trained and evaluated.
In Table˜6, we present the three-way routing results of SR-RAG and compare with always RAG as well as GenRead with the model before fine-tuning. Our main findings are as follows:
  • Adding PubMed as an additional knowledge source does not harm its effectiveness. The results on datasets other than PubMedQA aligns with the performance of two-source SR-RAG reported in the main text.
  • SR-RAG intelligently routes the best knowledge source for different workloads. For PubMedQA, PubHealth, and ARC, where PubMed can outperform Wiki by a large margin, SR-RAG frequently self-selects PubMedQA as its preferred knowledge source. On the other hand, on general factuality benchmarks such as PopQA and TriviaQA, SR-RAG selects Wikipedia more frequently. More importantly, SR-RAG achieves this behavior end-to-end without the need for any manual interventions.
Overall, this experiments confirms the effectiveness of extending SR-RAG to three sources and suggests a promising outlook for further extension.
@@ C.1 Experiment: Routing between General and Specialized Sources
Table 6: Three-way routing between retrieving from Wikipedia, retrieving from PubMed, or self-verbalizing knowledge. For each model we report task accuracy and the selection frequencies of the individual sources wiki/PubMed/self under SR-RAG’s LLM+kNN knowledge source routing scheme.
C.2 Discussion: Further Applications

### v3

@@ 3.4 Self-Routing RAG: Inference
As shown in Figures˜1 and 2, SR-RAG inference proceeds in a single left-to-right generation pass with three steps: source selection, knowledge collection, and answer generation.
@@
A common way to choose a source is to threshold the model probability of a source token, such as $p_{M}(\texttt{\textless Wiki\textgreater}\mid q)$ (DBLP:conf/iclr/AsaiWWSH24; repoformer). In practice, only relying on model’s uncertainty can be brittle and does not take into account model ability changes after the fine-tuning. To make source selection more robust and controllable, we augment token likelihoods with a dynamic kNN policy signal derived from additional rollouts. Concretely, we evaluate the fine-tuned model on a set of question-answer pairs [^3] Empirically, we reuse the training set so that no additional data or supervision is required.. For each pair, we compare the likelihood of generating the gold answer under different sources to decide a preferred source label. We then construct a policy datastore that maps each query to its preferred source, using the hidden representation at <EOQ> as the key. At test time, we retrieve the $k$ nearest neighbors of the query in this representation space and form a neighbor-induced distribution over sources $p_{D}(\texttt{\textless s\textgreater}\mid q)$ from label counts. Finally, we select the source by applying a threshold on the combined score:
$p_{M}(\texttt{\textless s\textgreater}\mid q)\times p_{D}(\texttt{\textless s\textgreater}\mid q).$ (11)
This combination adapts the decision boundary to the fine-tuned model’s behavior while preserving the efficiency of a single decoding pass. It also improves SR-RAG’s interpretability and controlability as the datastore consists of explicit source assignments, which can be audited, modified, or expanded by practitioners to steer routing behavior.
Knowledge Collection and Answer Generation
After selecting a source, SR-RAG gathers knowledge and generates the final answer within the same left-to-right pass. If the internal source $S_{i}$ is selected, the model uses greedy decoding to verbalize a single compact knowledge context, which serves as a cost-efficient approximation to multi-sample verbalization. If an external source $S_{e}$ is selected, decoding is paused to retrieve passages from the chosen index, which are appended as the knowledge context. In both cases, the system appends <EOK> and the model generates the final answer conditioned on the query and the knowledge.
@@ 4.1 Implementation Details of SR-RAG
The main experiments are performed on two knowledge sources [^4] We present multi-source extensions of SR-RAG in Appendix B.: the 2018 English Wikipedia (<Wiki>) as the external knowledge source, and the LLM itself (<Self>) as the internal knowledge source. We use the official Wikipedia embeddings released by karpukhin-etal-2020-dense and retrieve at a granularity of 100-word chunks. GenRead (DBLP:conf/iclr/0002IWXJ000023) is used to verbalize diverse knowledge contexts. GenRead clusters zero-shot knowledge verbalizations in as in-context demonstrations to verbalize diverse knowledge. We limit the verbalized knowledge chunks to a maximum of 150 tokens. From each knowledge source, we collect $n$ = 5 knowledge contexts.
@@
We fine-tune on a mixture of six short- and long-form knowledge-intensive datasets: Wizard of Wikipedia (DBLP:conf/iclr/DinanRSFAW19), Natural Questions (kwiatkowski-etal-2019-natural), FEVER (thorne-etal-2018-fever), OpenBookQA (mihaylov-etal-2018-suit), ARC-Easy (DBLP:journals/corr/abs-2102-03315), and ASQA (stelmakh-etal-2022-asqa). This mixture of 53,042 instances is a subset of the RAG instruction tuning data shown effective in DBLP:conf/iclr/AsaiWWSH24. After running the data construction algorithm, 46.9% of the instances are labeled with <Self> and the rest are labeled with <Wiki> as the preferred knowledge source, averaged across three LLMs. Full training details are provided in Section˜A.2.
@@ 4.2 Evaluation
We evaluate on a diverse set of four knowledge-intensive NLP tasks. PopQA (mallen-etal-2023-trust) is a free-form long-tail open-domain question answering dataset. Following DBLP:conf/iclr/AsaiWWSH24, we use the subset of 1,399 questions that aims to test knowledge of long-tail entities. TriviaQA (joshi-etal-2017-triviaqa) is an established open-domain question answering dataset that features relatively complex and diverse questions. We use the same test split and retrieval setup as in DBLP:conf/iclr/AsaiWWSH24. PubHealth (DBLP:journals/corr/abs-2304-03728) is a fact-checking dataset focusing on the public health domain. Finally, ARC Challenge (DBLP:journals/corr/abs-2102-03315) is a multiple-choice question answering dataset featuring grade-school level science questions. Following common practice, we perform lexical post-processing of the model’s output and report accuracy for PubHealth and ARC and substring matching for PopQA and TriviaQA.
@@
We compare SR-RAG with baselines that cover various training and inference strategies. (1) First, using the LLM before fine-tuning, we compare with either always retrieving or always verbalizing with GenRead. (2) As illustrated in Figure˜2, our main baseline is the state-of-the-art prior selective retrieval pipeline, combining the ideas from he-etal-2021-efficient, DBLP:conf/iclr/AsaiWWSH24, and repoformer. Specifically, the likelihoods of the LLM generating the answer with and without retrieval are used to create the knowledge source selection label. Then, we fine-tune the LLM for knowledge source selection (among $S_{e}$ and $\phi$) and generate the answer with optional retrieval. At inference, we apply a uniform threshold of 0.2 to the likelihood of the retrieval token for selective retrieval. (3) Always retrieving with the fine-tuned LLM after two-staged SR-RAG fine-tuning.
@@ 5.1 Knowledge Verbalization Alters Knowledge Source Preference
To motivate SR-RAG, we first show that knowledge verbalization can substantially change when an LLM benefits from retrieval. In a pilot study on Llama-2-7B-Chat using four datasets from our training mixture, we compare the likelihood of generating the gold answer under three conditions: no context (blue), the most helpful GenRead verbalization ($c_{i+}$, orange), and the most helpful retrieved passage ($c_{e+}$, red). As shown in Figure˜3, GenRead reverses the preferred knowledge source for a large fraction of instances, including 16% of Natural Questions and over 30% on the other datasets. This indicates that selective retrieval methods that omit verbalization can underestimate the LLM’s effective capability without retrieval, leading to suboptimal source preference labels and weaker performance when retrieval is skipped. Additional details of the study setup are provided in Section˜A.2.
@@
Table 2: Source selection accuracy. Llama-2-7B-Chat is used. SR-RAG achieves the best averaged performance in both evaluation settings.
Figure 4: Accuracy and system latency of SR-RAG. We show the performance of fine-tuned Llama-2-7B-Chat with different verbalization frequencies. SR-RAG’s source selection policy (marked with stars) achieves near-optimal accuracy-efficiency trade-off without dataset-specific thresholds.
@@ 5.2 SR-RAG Improves Accuracy with Reduced Retrieval Frequency
Table˜1 shows the end-to-end generation performance on three LLMs, demonstrating the advantage of SR-RAG over always retrieving, selective retrieval, and other baselines. Relative to always retrieving with the pretrained model, SR-RAG improves accuracy by 19.3% (Llama-2-7B-Chat), 13.4% (Phi-3.5-Mini-Instruct), and 5.9% (Qwen2.5-7B-Instruct). Compared to the selective retrieval baseline, SR-RAG achieves 8.5%/2.1%/4.7% higher accuracy while triggering 26%/40%/21% fewer retrievals across the three models, respectively. Notably, the selective retrieval baseline saves only 8–15% retrievals and achieves nearly identical accuracy to always retrieving, suggesting that skipping retrieval alone is not sufficient to reliably identify queries that can be handled without external context. In contrast, SR-RAG reduces retrieval by 20–40% while maintaining or improving accuracy, driven by accurate source selection and a stronger non-retrieval path.
Remarkably, without threshold tuning, SR-RAG can dynamically adapt its retrieval behavior. For PopQA, which emphasizes long-tail facts, SR-RAG retrieves external knowledge more frequently. For PubHealth and ARC, where the model can often rely on internal knowledge, SR-RAG skips retrieval more and can outperform always retrieving. We find that this flexibility also extends to selecting a specialized external source when multiple are present (Appendix˜B). We also present qualitative examples in the appendix (Figures˜7 and 8).
@@ 5.3 SR-RAG Makes Accurate Source Selection Decisions
Can SR-RAG choose the appropriate knowledge source? Table˜2 compares SR-RAG with Self-RAG (DBLP:conf/iclr/AsaiWWSH24) and an SR-RAG variant without kNN augmentation. We evaluate decision quality under two criteria: (1) abstaining from retrieval when skipping retrieval does not harm performance and (2) retrieving only when retrieval yields strictly better context. Under the first criterion, all methods perform well, and SR-RAG achieves the best overall accuracy. Under the stricter second criterion, SR-RAG substantially improves AUROC, outperforming Self-RAG by 14.7% on average. Removing kNN augmentation leads to large drops in both accuracy and AUROC, highlighting the benefit of adapting source selection to fine-tuning-induced shifts in model behavior. We visualize the hidden state space in Section˜C.2 to provide further intuitions.
@@ 5.4 System Efficiency
We measure end-to-end latency using Llama-2-7B-Chat under a batched inference setting (details in Section˜A.3). Figure˜4 shows the accuracy-latency trade-off as we vary retrieval frequency. As expected, latency improves as fewer instances invoke retrieval, and the optimal retrieval rate varies across datasets. Importantly, SR-RAG’s learned policy achieves a near-optimal accuracy-efficiency trade-off across datasets without dataset-specific threshold tuning. These results indicate that SR-RAG improves both accuracy and efficiency in a robust manner.
@@ 5.5 Ablation Study
We ablate SR-RAG components on Llama-2-7B-Chat in Table˜3. Disabling kNN-based routing (w/o. kNN) increases retrieval frequency and reduces accuracy, confirming that kNN augmentation improves robustness to ability shifts introduced by fine-tuning. Removing knowledge verbalization during preference labeling (w/o. kv. label) causes the model to over-rely on retrieval and degrades performance, showing that verbalization-aware supervision is important for accurate routing. Finally, ablating stage-2 preference alignment (training stage 2 with the stage-1 objective) reduces verbalization quality and leads to substantially more retrievals, consistent with weaker internal knowledge generation. Together, these results indicate that all three components contribute materially to SR-RAG’s overall gains. We provide additional analyses in Appendix˜C including the choice for source labeling heuristics, the size of the policy datastore, and other hyperparameters.
@@ 6 Conclusion
Appendix A SR-RAG: Further Details
A.1 List of Notations
@@
Table 4: A summary of the key symbols.
A.2 Training Details
@@
We implement GenRead following the setup in DBLP:conf/iclr/0002IWXJ000023. In the second verbalization round, five clusters with five in-context examples each are used. All datasets except ASQA follow the general prompt shown in Figure˜5. For ASQA, we include an additional instruction to handle ambiguity: “If the question is ambiguous, generate multiple documents for each possibility.
@@
During training, the loss is computed on the knowledge part only if <s> is <Self>. If <s> is <Wiki>, we augment $c_{e+}$ by randomly appending $\max(p-1,0)$ retrieved contexts, where $p$ is sampled from a Poisson distribution with $\lambda=2$. This data augmentation improves the LLM’s robustness to different retrieval strategies and various levels of retrieval quality. For stage 1 training, we use batch size 64, learning rate 1e-5, and fine-tune for 1 epoch. For stage 2 training, we use batch size 64, learning rate 5e-7, $\beta$ = 0.3 for DPO, and train for another epoch. All the experiments are performed on a machine with eight A800 (80GB) GPUs and a machine with eight A6000 GPUs. On eight A800 (80GB) GPUs, the two-stage training takes approximately 10 hours for a 7B-sized model.
A.3 Latency Formulation
@@
  • Verbalization Latency ($T_{v}$): The time for the LLM to verbalize parametric knowledge if the LLM itself is selected as the knowledge source.
@@
  • As the nearest-neighbor search only involves one matrix product for similarity calculation and one top-$k$ operation, $T_{d}$ is generally very small. In fact, we find $T_{d}$ (0.01s) <<<< $T_{g}$ (0.1s) << $T_{v}$ (1s). On the other hand, $T_{rs}$ is the major bottleneck of the pipeline. As a result, in an online setting, the system’s efficiency gain directly converges to the percentage of retrievals it is able to avoid.
@@
Table 6: Three-source results. We route between retrieving from Wikipedia, retrieving from PubMed, or self-verbalizing knowledge. For each model we report task accuracy and the selection frequencies of the individual sources wiki/PubMed/self under SR-RAG’s LLM+kNN knowledge source routing scheme.
Appendix B SR-RAG: Multi-Source Extensions
B.1 Experiment: Routing between General and Specialized Sources
We perform experiments on the PubMed corpus (ncbi-pubmed), which contains professional biomedical literature. Compared to Wikipedia and the LLM itself, PubMed has the unique advantage of featuring long-tail specialized biomedical knowledge. Meanwhile, it falls short at providing knowledge for many general factuality questions, posing a challenge to accurately routing the knowledge source.
Based on this new setup, we formulate SR-RAG with three knowledge source tokens: <Self>, <Wiki>, and <Pubmed>. We leverage the MedRAG toolkit (xiong-etal-2024-benchmarking) to implement retrieval on PubMed. The training and inference strategies of the two-source SR-RAG recipe naturally extends to this formulation and are followed in our experiments here. For training and evaluation, we additionally use PubMedQA (jin-etal-2019-pubmedqa), a widely used biomedical question answering dataset. PubMedQA features short questions extracted from PubMed abstracts and provides associated ground truth yes/no/maybe answers. We sample 10k artificially generated documents for training (combined with the data mixture used in the main text) and 500 examples from the expert-labeled subset for evaluation. The same set of LLMs from three families (Llama, Qwen, Phi) are trained and evaluated.
In Table˜6, we present the three-way routing results of SR-RAG and compare with always-RAG as well as GenRead with the model before fine-tuning. Our main findings are as follows:
  • Adding PubMed as an additional knowledge source does not harm its effectiveness. The results on datasets other than PubMedQA align with the performance of two-source SR-RAG reported in the main text.
  • SR-RAG intelligently routes the best knowledge source for different workloads. For PubMedQA, PubHealth, and ARC, where PubMed can outperform Wiki by a large margin, SR-RAG frequently self-selects PubMed as the knowledge source. On the other hand, on general factuality benchmarks such as PopQA and TriviaQA, SR-RAG selects Wikipedia more frequently. More importantly, SR-RAG achieves this behavior end-to-end without the need for any manual interventions.
Overall, these experiments confirm the effectiveness of extending SR-RAG to three sources and suggests a promising outlook for further extension.
B.2 Discussion: Further Applications